Find all solutions of the system (E) of linear equations

-27x+7y=14

-4x+y=15

Answer: (enter `1' for a true answer, and enter `0' for a false answer).

   The system (E) has infinitely many solutions
   The system (E) has exactly one solution
   The system (E) has no solutions

Moreover, these solutions are: t+span{a,b} for nonparallel vectors a and b.

For a vector that's not applicable, enter `-99' in each of its positions; if you mark a inapplicable, then b must be inapplicable also.
Answer #1

0 -27% + 74 -14 - unt y = 15 4 = 15+4% + © from equatim co and (2) -272+7(15 +420) = 14 -272 +105 +28 = 14 -272+284 = 14-105